A man is 320% yrs older than his son. Find the ratio of their ages
Answer by greenestamps(13200)



Choose an arbitrary age for the son that will make the calculations easier: suppose the son's age is 10. Then

320% of 10 = 10(3.20) = 32

320% OLDER THAN 10 = 10+32 = 42

The ratio of their ages is 10:42 = 5:21 = 1:4.2

If you understand percent increases, you don't have to go through all those calculations. An increase of 320% = 3.2 means the original number is multiplied by 1+3.2 = 4.2; the ratio is 1:4.2.